# Sharp error bounds for edge-element discretisations of the high-frequency Maxwell equations

### v2

@@ 2.3 The quantity $\gamma_{\rm dv}(P)$
Comparing (2.22) and (2.23), we see that, since the Galerkin error $u-u_{h}\not\in{\cal H}_{h}$, $u-u_{h}$ is not contained in the set of $w_{h}$ considered in (2.23). Nevertheless, controlling $\gamma_{\rm dv}(P)$ gives us a way to control $\Pi_{0}(u-u_{h})$, with Lemma 5.5 below showing that, for a certain projection $\Pi_{h}^{+}$, $\Pi_{0}\Pi^{+}_{h}(u-u_{h})$ is controlled by $\gamma_{\rm dv}(P)$, and then Lemma 5.1 controlling $\Pi_{0}(u-u_{h})$.
@@
In the Maxwell setting of Lemma 2.4, the $w_{h}$ considered in (2.23) are discretely $\epsilon$-divergence free. By Remark 2.6, if $\Pi_{1}w_{h}=w_{h}$ (i.e., $\Pi_{0}w_{h}=0$) then $w_{h}$ is $\epsilon$-divergence free. The quantity $\gamma_{\rm dv}(P)$ is therefore the familiar divergence conformity factor, measuring how much a finite-element function that is discretely $\epsilon$-divergence-free is not pointwise $\epsilon$-divergence-free, with this mismatch central to the analysis of the Maxwell FEM using Nédélec elements, as discussed in §1.2; see also [2, Lemma 5.2], [38, Lemma 4.5], [56, Lemma 7.6], and [10, Lemma 5.2] (with the notation $\gamma_{\rm dv}$ taken from [10, Equation 3.9]).
@@ 2.4 The main abstract theorem
Suppose that Assumption 2.1 holds for $P$ and that, for some $m\in\mathbb{Z}^{+}$ and spaces ${\cal Z}^{j},j=1,\ldots,m+1,$ Assumption 2.3 holds for both $P$ and $P^{*}$. Suppose that, given $f\in{\cal H}^{*}$, the solution to the equation $Pu=f$ exists and is unique.
Then there exist $C_{1},C_{2}$, and $C_{3}>0$ such that if
$\max\big{\{}\gamma_{\rm dv}(P),\gamma_{\rm dv}(P^{*})\big{\}}\leq C_{1}\,\,\text{ and }\,\,\Big{(}\left\|I-\Pi_{h}\right\|_{{\cal Z}^{m+1}\to{\cal H}}\Big{)}^{2}\big{(}1+\left\|(P^{*})^{-1}\Pi_{1}^{*}\right\|_{{\cal V}^{*}\to{\cal V}}\big{)}\leq C_{2},$ (2.24)
@@
$\left\|u-u_{h}\right\|_{{\cal H}}\leq C_{3}\Big{(}1+\left\|I-\Pi_{h}\right\|_{{\cal Z}^{m+1}\to{\cal H}}\big{(}1+\left\|(P^{*})^{-1}\Pi_{1}^{*}\right\|_{{\cal V}^{*}\to{\cal V}}\big{)}\Big{)}\left\|(I-\Pi_{h})u\right\|_{{\cal H}}.$ (2.25)
In addition, given $C_{\rm osc}>0$, there exists $C_{4}>0$ such that if $\|\iota^{-1}f\|_{{\cal Z}^{m+1}}\leq C_{\rm osc}\|f\|_{{\cal H}^{*}}$ and (2.24) holds, then
$\frac{\left\|u-u_{h}\right\|_{{\cal H}}}{\left\|u\right\|_{{\cal H}}}\leq C_{4}\Big{(}1+\left\|I-\Pi_{h}\right\|_{{\cal Z}^{m+1}\to{\cal H}}\big{(}1+\left\|(P^{*})^{-1}\Pi_{1}^{*}\right\|_{{\cal V}^{*}\to{\cal V}}\big{)}\Big{)}\left\|I-\Pi_{h}\right\|_{{\cal Z}^{m+1}\to{\cal H}}.$ (2.26)
@@
$\|\iota^{-1}\widetilde{\Pi}_{0}^{*}\iota\|_{{\cal Z}^{m-1}\to{\cal Z}^{m+1}}\leq C,$ (2.27)
@@
(i) By the order of quantifiers in Theorem 2.9, the constants $C_{1},C_{2},C_{3},$ and $C_{4}$ in the theorem depend on the quantities $\|P\|_{{\cal H}\to{\cal H}^{*}}$, $\|\mathcal{E}\|_{{\cal V}\to{\cal V}^{*}}$, $C_{\mathcal{E}}^{\prime}$, $C_{1}$, and $C_{2}$ in Assumption 2.1 and the quantity $C$ in Assumption 2.3.
@@ 3 Properties of $\Pi_{0}$ and $\Pi_{1}$
Now, from the matrix form of $P$ (4.2),
@@
By (4.3), the definition $\mathcal{E}_{10}^{*}:=(\Pi_{1}^{*}\mathcal{E}\Pi_{0})^{*}=\Pi_{0}^{*}\mathcal{E}^{*}\Pi_{1}$ (from (4.1)), (3.11), and (2.11) with $\mathsf{E}=\iota^{-1}\mathcal{E}^{*}$, for $j=2,\ldots,m+1$,
$\big{\|}\Pi_{0}u\big{\|}_{{\cal Z}^{j}}\leq C\big{\|}\iota^{-1}\Pi_{0}^{*}\mathcal{E}^{*}\Pi_{1}u\big{\|}_{{\cal Z}^{j}}=C\big{\|}\iota^{-1}\Pi_{0}^{*}\iota(\iota^{-1}\mathcal{E}^{*})\Pi_{1}u\big{\|}_{{\cal Z}^{j}}\leq C^{\prime}\big{\|}\Pi_{1}u\big{\|}_{{\cal Z}^{j}}.$ (4.9)
@@
$\left\|u\right\|_{{\cal Z}^{2}}\leq C\Big{(}1+\left\|(P^{*})^{-1}\Pi_{1}^{*}\right\|_{{\cal V}^{*}\to{\cal V}}\Big{)}\left\|g\right\|_{{\cal V}};$
the result then follows by combining this last displayed equation with (4.10). ∎
@@
From the matrix form of $P$ (4.2), $-\mathcal{E}_{00}\Pi_{0}u=\Pi_{0}^{*}f$. By (4.3), (3.11), and the bound on $f$ in (2.28),
@@
For the second assertion, now $-\mathcal{E}_{00}\Pi_{0}u=\widetilde{\Pi}_{0}^{*}f$. By (4.3) and the assumption (2.27),
@@
To prove the bound (2.25) in Theorem 2.9, it is sufficient to prove this bound under the assumption of existence. Indeed, by the assumed uniqueness of the solution to $Pu=f$, the bound (2.25) under the assumption of existence implies uniqueness of $u_{h}$, and uniqueness implies existence for the finite-dimensional Galerkin linear system. From now on, therefore, we assume that $u_{h}$ exists.
@@
The fact that $\Pi_{h}^{+}$ is well defined and bounded follows from Lemma 5.2 combined with the Lax–Milgram lemma [43], [49, Lemma 2.32], and the bound (5.4) then follows from Céa’s lemma [9], [18, Theorem 13.1]. The fact that $\Pi_{h}^{+}w_{h}=w_{h}$ for all $w_{h}\in{\cal H}_{h}$ follows from the facts that (i) $\Pi_{h}^{+}$ is well-defined, and (ii) if $w_{h}\in{\cal H}_{h}$, then $\Pi_{h}^{+}w_{h}=w_{h}$ is a solution of (5.3). ∎
@@ 6 Asymptotic quasi-optimality
$\big{(}1-C\gamma_{\rm dv}(P)\big{)}\left\|u-u_{h}\right\|_{{\cal H}}\leq C^{\prime}\left\|(I-\Pi_{h})u\right\|_{\cal H}+\left\|\Pi_{1}(u-u_{h})\right\|_{\cal H}.$ (6.3)
@@
$\qquad=\langle Pv,v\rangle_{{\cal H}^{*}\times{\cal H}}+\langle\mathcal{E}\Pi_{0}v,v\rangle_{{\cal V}^{*}\times{\cal V}}+\langle\mathcal{E}v,\Pi_{0}v\rangle_{{\cal V}^{*}\times{\cal V}}+\langle\mathcal{E}\Pi_{0}v,\Pi_{0}v\rangle_{{\cal V}^{*}\times{\cal V}}.$
@@
$\operatorname{Re}\big{\langle}P\Pi_{1}v,\Pi_{1}v\big{\rangle}_{{\cal H}^{*}\times{\cal H}}\leq\operatorname{Re}\big{\langle}Pv,v\big{\rangle}_{{\cal H}^{*}\times{\cal H}}+\varepsilon^{-1}\left\|\Pi_{0}v\right\|^{2}_{{\cal V}}+\varepsilon\left\|v\right\|^{2}_{{\cal V}}.$
@@ 7.2 Definition of $P^{\#}$ and $(P^{\#})^{-1}$
By definition, if $v\in\Pi_{1}{\cal H}$ then $\langle{\cal P}v,v\rangle_{{\cal H}^{*}\times{\cal H}}=\operatorname{Re}\langle Pv,v\rangle_{{\cal H}^{*}\times{\cal H}}$. Therefore, by (2.3),
$\big{\langle}{\cal P}v,v\big{\rangle}_{{\cal H}^{*}\times{\cal H}}\geq\left\|v\right\|^{2}_{{\cal H}}-(1+\|\mathcal{E}\|_{{\cal V}\to{\cal V}^{*}})\left\|v\right\|^{2}_{{\cal V}}\quad\text{ for all }v\in\Pi_{1}{\cal H}.$ (7.7)
@@
${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|v\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}:=\sqrt{Q(v,v)+(1+C)\left\|v\right\|_{V}^{2}}.$
@@
See, e.g., [63, Theorem VIII.15, Page 278], [37, Theorem 12.24, Page 360] (with [29] the original paper). ∎
@@
Furthermore $\eta^{-1}{\cal P}$ is a densely-defined, self-adjoint operator $\Pi_{1}{\cal V}\to\Pi_{1}{\cal V}$, with its spectrum bounded below by $-\|\mathcal{E}\|_{{\cal V}\to{\cal V}^{*}}$.
@@
We now check the assumptions of Theorem 7.6. Since ${\cal H}$ is dense in ${\cal V}$, and $\Pi_{1}:{\cal V}\to{\cal V}$ is bounded, $H=\Pi_{1}{\cal H}$ is dense in $V=\Pi_{1}{\cal V}$. By its definition (7.6), ${\cal P}:\Pi_{1}{\cal H}\to\Pi_{1}^{*}{\cal H}^{*}$ is self adjoint; i.e., Assumption (i) of Theorem 7.6 is satisifed. The Gårding inequality (7.7) then implies that Assumptions (ii) and (iii) of Theorem 7.6 are satisfied with $C=C_{\mathcal{E}}$.
@@
For all $\varepsilon>0$, $\eta^{-1}{\cal P}+\|\mathcal{E}\|_{{\cal V}\to{\cal V}^{*}}+\varepsilon:{\rm Dom}(\eta^{-1}{\cal P})\to\Pi_{1}{\cal V}$ is then invertible by a variant of the Lax–Milgram lemma for densely-defined operators; see, e.g., [37, Theorem 12.18] or the proof of [63, Theorem VIII.15]. Thus the spectrum of $\eta^{-1}{\cal P}$ (i.e., the set of $\lambda$ such that $\eta^{-1}{\cal P}-\lambda:{\rm Dom}(\eta^{-1}{\cal P})\to\Pi_{1}{\cal V}$ is not invertible) is bounded below by $-\|\mathcal{E}\|_{{\cal V}\to{\cal V}^{*}}$. ∎
@@
Let ${\cal L}$ be a densely-defined, self-adjoint operator on a Hilbert space $V$, and let $\sigma({\cal L})$ denote its spectrum.
@@
We now use the Gårding inequality (7.7) to replace the ${\cal V}$ norm on the right-hand side of (7.12) by a ${\cal H}$ norm and obtain (7.11). Since $\psi^{2}\geq 0$, $S^{2}\geq 0$. Using this, along with (7.7) and (7.12), we find that, for all $\varepsilon>0$ and $v\in\Pi_{1}{\cal H}$,

### v3

@@ 2.3 The quantity $\gamma_{\rm dv}(P)$
Comparing (2.22) and (2.23), we see that, since the Galerkin error $u_{h}\not\in{\cal H}_{h}, uu$$u_{h}$ is not contained in the set of $w_{h}$ considered in (2.23). Nevertheless, controlling $\gamma_{\rm dv}(P)$ gives us a way to control $\Pi_{0}(u-u_{h})$, with Lemma 5.5 below showing that, for a certain projection $\Pi_{h}^{+}$, $\Pi_{0}\Pi^{+}_{h}(u-u_{h})$ is controlled by $\gamma_{\rm dv}(P)$, and then Lemma 5.1 controlling $\Pi_{0}(u-u_{h})$.
@@
In the Maxwell setting of Lemma 2.4, the $w_{h}$ considered in (2.23) are discretely $\epsilon$-divergence free. By Remark 2.6, if $\Pi_{1}w_{h}=w_{h}$ (i.e., $\Pi_{0}w_{h}=0$) then $w_{h}$ is $\epsilon$-divergence free. The quantity $\gamma_{\rm dv}(P)$ is therefore the familiar divergence conformity factor, measuring how much a finite-element function that is discretely $\epsilon$-divergence-free is not pointwise $\epsilon$-divergence-free, with this mismatch central to the analysis of the Maxwell FEM using Nédélec elements, as discussed in §1.2; see also [2, Lemma 5.2], [38, Lemma 4.5], [57, Lemma 7.6], and [10, Lemma 5.2] (with the notation $\gamma_{\rm dv}$ taken from [10, Equation 3.9]).
@@ 2.4 The main abstract theorem
Fix the constants in Assumptions 2.1 and 2.3 (i.e., $C_{{\mathcal{E}}}^{\prime},C_{1},C_{2}$ in Assumption 2.1 and $C_{\rm reg}$} in Assumption 2.3), fix $m\in\mathbb{Z}^{+}$ and spaces ${\cal Z}^{j},j=1,\ldots,m+1$, and let $C_{\rm osc},C_{\Pi_{0}}>0$. Then there exist $c,C>0$ such that for all $P$ such that
  • Assumption 2.1 holds for $P$ (with $C_{{\mathcal{E}}}^{\prime},C_{1},C_{2}$ fixed above),
  • Assumption 2.3 holds for both $P$ and $P^{*}$ (with $C_{{\mathcal{E}}}^{\prime},C_{1},C_{2},{C_{\rm reg}}$, $m$, and ${\cal Z}^{j},j=1,\ldots,m+1$ fixed above) and,
  • given $f\in{\cal H}^{*}$, the solution to the equation $Pu=f$ exists and is unique,
if
$\gamma_{\rm dv}(P)+\gamma_{\rm dv}(P^{*})+\Big(\left\|I-\Pi_{h}\right\|_{{\cal Z}^{m+1}\to{\cal H}}\Big)^{2}\big(1+\left\|(P^{*})^{-1}\Pi_{1}^{*}\right\|_{{\cal V}^{*}\to{\cal V}}\big)\leq c,$ (2.24)
@@
$\left\|u-u_{h}\right\|_{{\cal H}}\leq C\Big(1+\left\|I-\Pi_{h}\right\|_{{\cal Z}^{m+1}\to{\cal H}}\big(1+\left\|(P^{*})^{-1}\Pi_{1}^{*}\right\|_{{\cal V}^{*}\to{\cal V}}\big)\Big)\left\|(I-\Pi_{h})u\right\|_{{\cal H}}.$ (2.25)
In addition, if $\|\iota^{-1}f\|_{{\cal Z}^{m+1}}\leq C_{\rm osc}\|f\|_{{\cal H}^{*}}$ and (2.24) holds, then
$\frac{\left\|u-u_{h}\right\|_{{\cal H}}}{\left\|u\right\|_{{\cal H}}}\leq C\Big(1+\left\|I-\Pi_{h}\right\|_{{\cal Z}^{m+1}\to{\cal H}}\big(1+\left\|(P^{*})^{-1}\Pi_{1}^{*}\right\|_{{\cal V}^{*}\to{\cal V}}\big)\Big)\left\|I-\Pi_{h}\right\|_{{\cal Z}^{m+1}\to{\cal H}}.$ (2.26)
@@
$\|\iota^{-1}\widetilde{\Pi}_{0}^{*}\iota\|_{{\cal Z}^{m-1}\to{\cal Z}^{m+1}}\leq C_{\Pi_{0}},$ (2.27)
@@
(i) By the order of quantifiers in Theorem 2.9, $c$ and $C$ in the theorem depend on $C_{\mathcal{E}}^{\prime}$, $C_{1}$, and $C_{2}$ in Assumption 2.1, $C_{\rm reg}$} in Assumption 2.3, and the fixed constants $C_{\rm osc}$ and $C_{\Pi_{0}}$. In the proof of Theorem 2.9, $C_{\mathcal{E}}^{\prime}$, $C_{1}$, $C_{2}$, $C_{\rm reg}$}, $C_{\rm osc}$ and $C_{\Pi_{0}}$ are fixed at the outset, but, for brevity, this is not stated explicitly in each intermediate result. In the proofs, the letter $C$ denotes a constant that, in principle, depends on $C_{\mathcal{E}}^{\prime}$, $C_{1}$, $C_{2}$, $C_{\rm reg}$}, $C_{\rm osc}$ and $C_{\Pi_{0}}$, but nothing else.
@@ 3 Properties of $\Pi_{0}$ and $\Pi_{1}$
Now, from the matrix form of $P$ (4.2) and the equation $P^{*}u=\Pi_{1}^{*}\iota g$,
@@
By the combination of (4.3), (4.8), the definition $\mathcal{E}_{10}^{*}:=(\Pi_{1}^{*}\mathcal{E}\Pi_{0})^{*}=\Pi_{0}^{*}\mathcal{E}^{*}\Pi_{1}$ (from (4.1)), (3.11), and (2.11) with $\mathsf{E}=\iota^{-1}\mathcal{E}^{*}$, for $j=2,\ldots,m+1$,
$\left\|\Pi_{0}u\right\|_{{\cal Z}^{j}}$ $\leq C\left\|\iota^{-1}\mathcal{E}_{00}^{*}\Pi_{0}u\right\|_{{\cal Z}^{j}}$
$=C\left\|\iota^{-1}\mathcal{E}_{10}^{*}\Pi_{1}u\right\|_{{\cal Z}^{j}}$
$\leq C\big\|\iota^{-1}\Pi_{0}^{*}\mathcal{E}^{*}\Pi_{1}u\big\|_{{\cal Z}^{j}}=C\big\|\iota^{-1}\Pi_{0}^{*}\iota(\iota^{-1}\mathcal{E}^{*})\Pi_{1}u\big\|_{{\cal Z}^{j}}\leq C^{\prime}\big\|\Pi_{1}u\big\|_{{\cal Z}^{j}}.$ (4.9)
@@
$\left\|u\right\|_{{\cal Z}^{2}}\leq C\Big(1+\left\|(P^{*})^{-1}\Pi_{1}^{*}\right\|_{{\cal V}^{*}\to{\cal V}}\Big)\left\|g\right\|_{{\cal V}}.$ (4.11)
The result then follows by the combination of (4.10), (4.11), and induction. ∎
@@
From the matrix form of $P$ (4.2), $\mathcal{E}_{00}\Pi_{0}u=\Pi_{0}^{*}f$. By (4.3), (3.11), and the bound on $f$ in (2.28),
@@
For the second assertion, now $\mathcal{E}_{00}\Pi_{0}u=\widetilde{\Pi}_{0}^{*}f$. By (4.3) and the assumption (2.27),
@@
To prove the bound (2.25) in Theorem 2.9, we claim that it is sufficient to prove this bound under the assumption that $u_{h}$ exists. To justify this claim, observe that since $u_{h}$ is the solution of a finite-dimensional linear system, the statements “under the assumption that $u_{h}$ exists, $u_{h}$ is unique” and $u_{h}$ exists and is unique” are equivalent. Once the bound (2.25) is established under the assumption that $u_{h}$ exists, setting $f=0$ and using that then (by one of the assumptions in Theorem 2.9) $u=0$, we find that $u_{h}=0$; i.e., under the assumption that $u_{h}$ exists, $u_{h}$ is unique, and hence $u_{h}$ exists and is unique by the above equivalence. From now on, therefore, we assume that $u_{h}$ exists and seek to prove (2.25).
@@
The fact that $\Pi_{h}^{+}$ is well defined and bounded follows from Lemma 5.2 combined with the Lax–Milgram lemma [43], [50, Lemma 2.32], and the bound (5.4) then follows from Céa’s lemma [9], [18, Theorem 13.1]. The fact that $\Pi_{h}^{+}w_{h}=w_{h}$ for all $w_{h}\in{\cal H}_{h}$ follows from the facts that (i) $\Pi_{h}^{+}$ is well-defined, and (ii) if $w_{h}\in{\cal H}_{h}$, then $\Pi_{h}^{+}w_{h}=w_{h}$ is a solution of (5.3). ∎
@@ 6 Asymptotic quasi-optimality
$\big(1-C\gamma_{\rm dv}(P)\big)\left\|u-u_{h}\right\|_{{\cal H}}\leq C\left\|(I-\Pi_{h})u\right\|_{\cal H}+\left\|\Pi_{1}(u-u_{h})\right\|_{\cal H}.$ (6.3)
@@
$\qquad=\langle Pv,v\rangle_{{\cal H}^{*}\times{\cal H}}+\langle\mathcal{E}\Pi_{0}v,v\rangle_{{\cal V}^{*}\times{\cal V}}+\langle\mathcal{E}v,\Pi_{0}v\rangle_{{\cal V}^{*}\times{\cal V}}-\langle\mathcal{E}\Pi_{0}v,\Pi_{0}v\rangle_{{\cal V}^{*}\times{\cal V}}.$
@@
$\operatorname{Re}\big\langle P\Pi_{1}v,\Pi_{1}v\big\rangle_{{\cal H}^{*}\times{\cal H}}\leq\operatorname{Re}\big\langle Pv,v\big\rangle_{{\cal H}^{*}\times{\cal H}}+C\big(\varepsilon^{-1}\left\|\Pi_{0}v\right\|^{2}_{{\cal V}}+\varepsilon\left\|v\right\|^{2}_{{\cal V}}\big).$
@@ 7.2 Definition of $P^{\#}$ and $(P^{\#})^{-1}$
By definition, if $v\in\Pi_{1}{\cal H}$ then $\langle{\cal P}v,v\rangle_{\Pi_{1}^{*}{\cal H}^{*}\times\Pi_{1}{\cal H}}=\operatorname{Re}\langle Pv,v\rangle_{{\cal H}^{*}\times{\cal H}}$. Therefore, by (2.3),
$\big\langle{\cal P}v,v\big\rangle_{\Pi_{1}^{*}{\cal H}^{*}\times\Pi_{1}{\cal H}}\geq\left\|v\right\|^{2}_{{\cal H}}-(1+\|\mathcal{E}\|_{{\cal V}\to{\cal V}^{*}})\left\|v\right\|^{2}_{{\cal V}}\quad\text{ for all }v\in\Pi_{1}{\cal H}.$ (7.7)
@@
[…]
@@
See, e.g., [64, Theorem VIII.15, Page 278], [37, Theorem 12.24, Page 360] (with [29] the original paper). ∎
@@
Furthermore $\eta^{-1}{\cal P}$ is a densely-defined, self-adjoint operator $\Pi_{1}{\cal V}\to\Pi_{1}{\cal V}$, with its spectrum bounded below by $\|\mathcal{E}\|_{{\cal V}\to{\cal V}^{*}}$.
@@
[…]
@@
For all $\varepsilon>0$, $\eta^{-1}{\cal P}+\|\mathcal{E}\|_{{\cal V}\to{\cal V}^{*}}+\varepsilon:{\rm Dom}(\eta^{-1}{\cal P})\to\Pi_{1}{\cal V}$ is then invertible by a variant of the Lax–Milgram lemma for densely-defined operators; see, e.g., [37, Theorem 12.18] or the proof of [64, Theorem VIII.15]. Thus the spectrum of $\eta^{-1}{\cal P}$ (i.e., the set of $\lambda$ such that $\eta^{-1}{\cal P}-\lambda:{\rm Dom}(\eta^{-1}{\cal P})\to\Pi_{1}{\cal V}$ is not invertible) is bounded below by $\|\mathcal{E}\|_{{\cal V}\to{\cal V}^{*}}$. ∎
@@
[…]
@@
[…]